# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v2

@@ C.3 Reward Function Discussion
C.4 Computing Infrastructure
@@ Appendix D Omitted Experiments
D.1 Visualization
Figure 6: The $(s,a,s^{\prime})$ distributions generated by SAMBO result in the walker2d-medium-expert task in D4RL benchmark. Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
@@ D.1 Visualization
D.2 Detailed Ablation Discussion
SAMBO is regulated by the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
The model bias adjustment coefficient $\alpha$ controls the degree of adjustment to model-generated data. A larger $\alpha$ leads to excessive conservatism, restricting exploration even in well-performing regions and potentially causing suboptimal performance. Conversely, a smaller $\alpha$ may encourage excessive optimism in exploring risky regions, impairing performance. To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task. The results, shown in Fig. 7 (a), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance. Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration. This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy. A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy. However, this also introduces a heightened risk of taking unsafe actions. As shown in Fig. 7 (b), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness. A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability. Thus, $\beta$ should also be set appropriately to balance exploration and stability.
(a) Tuning for $\alpha$
(b) Tuning for $\beta$
Figure 7: Illustration of hyperparameter experiments in the HalfCheetah-random task in D4RL benchmark.
D.3 Performance Compare
We also compared the performance curves of model-based approaches on other tasks in the NeoRL benchmark. The curve for MOBILE was obtained by executing the official implementation. Since MOPO does not have an official implementation on NeoRL, we report its score based on the results presented in the original NeoRL paper.
As shown in Figures 10 to 16, SAMBO outperforms in all Hopper and Walker2d tasks, both of which are more complex and challenging. Although it exhibits slower convergence initially in the Walker2d-M and Walker2d-H tasks, its performance improves rapidly as the experiment progresses, ultimately achieving state-of-the-art (SOTA) levels. On the other hand, SAMBO also demonstrates comparable performance in simpler HalfCheetah tasks.
Figure 8: Hopper-L
Figure 9: Hopper-M
Figure 10: Hopper-H
Figure 11: Walker2d-L
Figure 12: Walker2d-M
Figure 13: Walker2d-H
Figure 14: HalfCheetah-L
Figure 15: HalfCheetah-M
Figure 16: HalfCheetah-H

### v3

@@ C.3. Reward Function Discussion
    In this case, $\epsilon(G;s)=0,\forall s\in\mathcal{S}$ in Theorem 4, but this derived surrogate objective is not a lower bound of the original RL objective.
C.4. Computing Infrastructure
@@ Appendix D Omitted Experiments
D.1. Visualization
Figure 6. The $(s,a,s^{\prime})$ distributions generated by SAMBO in the walker2d-medium-expert task in D4RL benchmark. Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
@@ D.1. Visualization
Figure 7. Hopper-L
Figure 8. Hopper-M
Figure 9. Hopper-H
Figure 10. Walker2d-L
Figure 11. Walker2d-M
Figure 12. Walker2d-H
Figure 13. HalfCheetah-L
Figure 14. HalfCheetah-M
Figure 15. HalfCheetah-H
D.2. Performance Comparison
We also compared the performance curves of model-based approaches on other tasks in the NeoRL benchmark. The curve for MOBILE was obtained by executing the official implementation. Since MOPO does not have an official implementation on NeoRL, we report its score based on the results presented in the original NeoRL paper. As shown in Figures 15 to 15, SAMBO outperforms in all Hopper and Walker2d tasks, both of which are more complex and challenging. Although it exhibits slower convergence initially in the Walker2d-M and Walker2d-H tasks, its performance improves rapidly as the experiment progresses, ultimately achieving SOTA levels. On the other hand, SAMBO does not achieve strong performance on the HalfCheetah tasks. We attribute this to the fact that the HalfCheetah score mainly depends on achieving high running speed rather than mere stability. Since SAMBO’s dual conservatism with respect to model bias and policy shift places greater emphasis on stability, it performs less competitively on HalfCheetah, while demonstrating strong results on Hopper and Walker2d, where stability is more critical.